# Twisted Superfields

### v7

@@ 3.2 Observed boson masses
The Z boson mass in this model is also primarily determined by $\bar{\phi}_{21}^{2}$. Due to the form of eq (2.16), the Z boson mass in this model differs slightly from the Standard Model expression. Namely, the mass of the Z boson is:
@@
where the angles are defined by coupling constants as in eq (2.17). In section 5 when discussing the $Z$ boson width, a scenario is presented where $\sin\phi_{Z}\simeq 0.1$ at the $Z$ boson scale (getting much smaller at the $Z^{\prime}$ boson scale). The implications for precision $W$ mass measurements are discussed in that section. In most scenarios, $\sin\phi_{Z}$ should not be larger than that (and possibly much smaller), so the following approximation can be used for most purposes in this paper:
@@
where it has been assumed that $g_{+}=g$ at the unification scale and that $g_{Y^{\prime}}$ is negligible due to reasons discussed below. The approximation has also assumed that $\bar{\phi}^{2}_{21}\gg\bar{\tilde{\phi}}_{22},\bar{\varphi}_{2}$.
@@
where the scalar mass $m_{21}$ is from eq (1.18). The fourth line above comes from the relations in eqs (2.17) and (3.6) along with the approximation that $g_{U}\simeq g_{2}$. In the $Z$ width part of section 5, it is suggested that $m_{21}$ may be very small. In that case, roughly 28 GeV of the Higgs Boson mass in this model is assumed to come from radiative corrections.
@@
On the other hand, interactions of $h_{21}$ with other quarks or leptons are different in this model than in the Standard Model. As described later in this section, leptons and quarks (other than the top) acquire masses from different mechanisms, not from Yukawa interactions with the Higgs Boson. Those particles do have interactions with $h_{21}$ via the nonperturbative terms described in the Appendix. More work would need to be done to determine whether those interactions were consistent with observations.
@@ 3.3 Quark masses and mixing
$\mathit{M}_{u}=\left(\begin{array}[]{ccc}{\hat{g}\bar{\tilde{\phi}}_{21}}&{\tfrac{1}{\sqrt{2}}\tilde{\Gamma}_{211}\bar{\phi}_{21}}&{\tilde{m}_{211}}\\ {\tilde{M}_{G}^{(1)}-\Delta}&{\tilde{m}_{\Phi}}&{\tfrac{1}{\sqrt{2}}\bar{\tilde{\phi}}_{21}\tilde{\Gamma}_{211}}\\ {\tilde{m}_{\lambda}}&{\tilde{M}_{G}^{(1)}+\Delta}&{\hat{g}\bar{\phi}_{21}}\end{array}\right),$
where the gaugino coupling $\hat{g}$ is discussed below in eq (3.14). A tilde is put on tree-level masses and superpotential couplings to show that they include quantum modifications from the appendix. For all of the non-neutrino fermion mass matrices in this paper (including the above matrix) the rows have an upper undotted spin index (using the convention of [7]) while the columns have a lower undotted spin index (e.g. $\tilde{u}^{\left(1\right)\alpha}u_{W\alpha}^{G}$). Those indices are summed over, and they are suppressed.
@@
The first- and second-generation quark interactions with the W boson, however, differ from those of the Standard Model. For the second-generation up-type quark (mostly charm), both its left- and right-chiral components interact with the W boson. For the first-generation up-type quark (mostly up), only its right-chiral component interacts with the W boson. Nuclear beta decay experiments are consistent with the W boson connecting left-chiral protons to left-chiral neutrons. If the spin of a proton was primarily just the sum of the spins of its three valence quarks, then the right-chiral W boson interactions of this model’s up (and down) quarks might not be consistent with nuclear beta decay experiments. However, the spin of the proton mostly comes from gluons and orbital angular momentum [20], so this model’s right-chiral interactions are not inconsistent with nuclear beta decay data.
@@
Instead, the following approach is used: Fits are made below to quark masses and CKM data. Those fits produce a numerical value for $\hat{g}\bar{\phi}_{21}$. By comparing that numerical value to the mass of the W boson given by $\sqrt{2}M_{W}\simeq g_{2}\bar{\phi}_{21}$, the following ratio can be extracted:
$\hat{g}\simeq 1.5g_{2}\simeq 0.8g_{3},$ (3.14)
@@
An interesting comparison is provided by one of the results from the Minimal Supersymmetric Standard Model [21, 22, 23]. In the MSSM, the Yukawa coupling for the top quark can be approximated by [11]
$\lambda_{t}^{2}\left(M_{Z}\right)$ $=\frac{7g_{3}^{2}\left(M_{Z}\right)}{18}\left(1-\left(\frac{\alpha_{3}\left(M_{U}\right)}{\alpha_{3}\left(M_{Z}\right)}\right)^{{7\mathord{\left/{\vphantom{79}}\right.\kern-1.2pt}9}}\right)^{-1}$ (3.15)
$\sim\left(0.7g_{3}\right)^{2},$
where $M_{U}$ is the unification scale (and 7/18 is called the Pendleton Ross fixed point). The value of the strong coupling at the unification scale in the above formula is taken from section 4 of this paper.
Before wrapping up this discussion of up-type quarks, it is worth repeating that in this model, the one vev $\bar{\phi}_{21}$ chosen by the vacuum is presumed to determine not only the masses of the W, Z and Higgs bosons, but also the mass of the top quark without the addition of a separate Yukawa coupling.
@@
$\rm{Rows:}\,\tilde{d}^{A},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{\left(2\right)}\,\,\,\rm{Columns:}\,d_{W}^{G},d^{\left(2\right)},d^{A},d_{W}^{A}$ (3.16)
$M_{d}=\left(\begin{array}[]{cccc}{...}&{...}&{\tilde{m}^{\prime\prime}_{\Phi}}&{\hat{g}\bar{\varphi}_{2}}\\ {...}&{\tfrac{1}{\sqrt{2}}\bar{\tilde{\phi}}_{22}\tilde{\Gamma}_{222}}&{\tfrac{1}{\sqrt{2}}\bar{\varphi}_{2}\tilde{\Gamma}_{\Phi}}&{\tilde{m}^{\prime}_{\Phi}}\\ {\tilde{m}^{\prime}_{\lambda}}&{...}&{...}&{...}\\ {\hat{g}\bar{\tilde{\phi}}_{22}}&{\tilde{m}_{222}}&{...}&{...}\end{array}\right),$
@@
Construction of the CKM matrix in this model is more involved than in the Standard Model. As in the Standard Model, the first step is to find unitary matrices $V_{L}^{U}$, $V_{R}^{U}$, $V_{L}^{D}$, and $V_{R}^{D}$ that diagonalize the quark mass matrices:
$V_{L}^{U}\mathit{M}_{u}V_{R}^{U{\dagger}}\,\,\rm{and}$ (3.17)
$V_{L}^{D}\mathit{M}_{d}V_{R}^{D{\dagger}}\,\,\textrm{are diagonal.}$.
Unlike in the Standard Model where only the $V_{L}^{U,D}$ matrices have connections to the W boson, in this model, all four of the matrices have connections to the W.
@@
The placement of the 1’s in the above matrices is based on which quark fields in eqs (3.12) and (3.16) have a $W$ subscript (signifying that they interact with the W boson to change an up-type quark to a down-type quark and vice versa). In this model the “CKM” matrix is not a square, it is 3x4.
Also, in this model, a different CKM matrix should be used depending on whether a vector current or axial vector current process is being considered. Specifically:
@@
To fit the parameters in the up- and down-type quark mass matrices, it is first assumed that the fourth down-type quark has a mass of 2.8 GeV. The reasons and experimental justifications for this choice are discussed in section 5.
A fit of this model’s parameters was performed using the six known quark masses, this new quark mass, and measured CKM data, both for vector-current and axial-vector-current decays. The following values (in GeV) were obtained for elements of the mass matrices:
$\mathit{M}_{u}=\left(\begin{array}[]{ccc}{0.00131}&{-0.00191}&{0.0122}\\ {0.6493}&{1.1126}&{-5.3477}\\ {1.7379}&{-1.0843}&{171.94}\end{array}\right).$
$\rm{Rows:}\,\tilde{d}^{A},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{\left(2\right)}\,\,\,\rm{Columns:}\,d_{W}^{G},d^{\left(2\right)},d^{A},d_{W}^{A}$ (3.21)
$\mathit{M}_{d}=\left(\begin{array}[]{cccc}{-0.0338}&{0.0137}&{0.0246}&{-0.0527}\\ {-0.0252}&{-0.0725}&{-0.0636}&{-0.0804}\\ {0.1054}&{1.4894}&{-3.8554}&{0.1329}\\ {1.4717}&{-0.0800}&{0.2408}&{2.3052}\end{array}\right).$
To find the unitary matrices that diagonalized the above matrices and generated CKM matrices, each mass matrix was first multipled by its transpose, then matrices of left- and right- eigenvectors were found. The allowed flexibility to introduce complex phases was not used for the fits of this paper, so no attempt was made to fit the experimentally measured complex phases of the CKM matrix.
With the above mass matrices, the model generates the following for the quark masses and the CKM matrices:
$m_{u}=0.002,\,\,\,m_{c}=1.28,\,\,\,m_{t}=172$ (3.22)
$m_{d}=0.005,\,\,\,m_{s}=0.095,\,\,\,m_{b}=4.18,\,\,\,m_{f}=2.8$
$\left|V_{\rm{CKM}}^{+}\right|=\left(\begin{array}[]{cccc}{0.9745}&{0.2239}&{0.0043}&{0.0094}\\ {0.2323}&{0.9661}&{0.0388}&{0.9720}\\ {0.0088}&{0.0402}&{0.9973}&{0.0619}\end{array}\right)$
$\left|V_{\rm{CKM}}^{-}\right|=\left(\begin{array}[]{cccc}{0.9742}&{0.2252}&{0.0042}&{0.0093}\\ {0.2165}&{0.9805}&{0.0394}&{1.0278}\\ {0.0144}&{0.0349}&{0.9974}&{0.0621}\end{array}\right),$
where all mass values are in GeV, and the subscript $f$ was used to denote the fourth down-type quark.
@@
3.4 Charged leptons
After unification scale symmetry breaking but before electroweak symmetry breaking, this model has the following lepton content: (i) three fundamental-representation isodoublet pairs, (ii) three conjugate-representation isodoublet pairs, (iii) a fundamental and a conjugate isosinglet charged lepton, and (iv) 2 fundamental and 2 conjugate isosinglet neutrinos. Isosinglet leptons do not interact with the W boson.
This section maps this model’s charged leptons to the observed leptons. The first part of this section maps the flavor 1 charged leptons to the electron and an as-yet-undiscovered heavy lepton. The second part maps the flavor 2 and 3 charged leptons to the muon and tau lepton.
From eqs (A.10), (A.16) and (A.17) of the appendix, the flavor-1 lepton mass matrix (for Dirac masses) takes the following form:
$\begin{array}[]{c}{\left(\begin{array}[]{ccc}{\tilde{\nu}_{W}^{\left(1\right)}}&{\tilde{e}_{W}^{\left(1\right)}}&{\tilde{e}^{\left(1\right)}}\end{array}\right)}\\ {}\hfil\\ {}\hfil\end{array}\left(\begin{array}[]{ccc}{\tilde{m}_{111}}&{0}&{0}\\ {0}&{\tilde{m}^{\prime}_{111}}&{M_{\Omega}}\\ {0}&{m_{e}}&{\tilde{m}^{\prime}_{111}}\end{array}\right)\left(\begin{array}[]{c}{\nu_{W}^{\left(1\right)}}\\ {e_{W}^{\left(1\right)}}\\ {e^{\left(1\right)}}\end{array}\right)\,,$ (3.23)
where $m_{e}$ is the mass of the electron and $M_{\Omega}$ is the mass of a new, proposed heavy lepton. This mass should be in the range 103 GeV $<M_{\Omega}<2\times 10^{6}$ GeV. The lower limit comes from current 95% confidence exclusion limit from heavy lepton searches [26]. The upper limit is from the anomaly-based analysis surrounding eq (4.1).
From eq (2.19), the last three diagonal components of the photon field when acting on flavor-1 fundamental-representation fields are $e\left(0,1,1\right)$. As a result, if the left- and right-chiral electron components are defined via $e_{L}^{\dagger}=e_{W}^{\left(1\right)}$ and $e_{R}=\tilde{e}^{\left(1\right)}$, then those fields have the correct electric charges of $+1$ and $-1$ as well as the correct isodoublet and isosinglet designations. In the limit that $\tilde{m}^{\prime}_{111}\ll m_{\Omega}$, the flavor 1 leptons can therefore be identified as follows:
$\begin{array}[]{ccc}{\nu_{W}^{(1)}=\nu_{eL}^{\dagger}}&{e_{W}^{(1)}=e_{L}^{\dagger}}&{e^{(1)}=\Omega_{L}^{\dagger}}\\ {\tilde{\nu}_{W}^{(1)}=\nu_{\Omega R}}&{\tilde{e}_{W}^{(1)}=\Omega_{R}}&{\tilde{e}^{(1)}=e_{R},}\end{array}$ (3.24)
where $\Omega$ is the hypothesized new heavy lepton. From this identification, it can be seen that the right-chiral component (not left-chiral) of the negatively charged Omega lepton is the one that interacts with the W boson, forming an isodoublet with its right-chiral neutrino.
The electron not only has the same interactions with the photon and W boson as in the Standard Model, it also has the same interactions with the Z boson, as can be seen from eqs (2.18) and (2.20).
In this model, the muon and tau lepton come from the flavor 2/3 lepton sector of the theory. From eqs (A.10), (A.16) and (A.17) of the appendix, the flavor 2/3 lepton mass matrix takes the form:

### v8

@@ 3.2 Observed boson masses
The $Z$ boson mass in this model is also primarily determined by $\bar{\phi}_{21}^{2}$. Due to the form of eq (2.16), the $Z$ boson mass in this model differs slightly from the Standard Model expression. The mass of the $Z$ boson is:
@@
where the angles are defined by coupling constants as in eq (2.17). In section 5 when discussing the $Z^{\prime}$ boson, $\sin\phi_{Z}$ is estimated to be on the order of $10^{-3}$ at the scale of $\sim$17 MeV. It could be a little larger at the mass of the $Z$ boson, but should still be very small. For that reason, the following approximation can be used for most purposes in this paper:
@@
where it has been assumed that $g_{+}=g$ at the unification scale and that $g_{Y^{\prime}}$ is negligible due to reasons discussed below. The approximation has also assumed that $\bar{\phi}^{2}_{21}\gg\bar{\phi}_{22},\bar{\varphi}_{2}$.
@@
where the scalar mass $m_{21}$ is from eq (1.18). The third line above comes from the relations in eqs (2.17) and (3.6) along with the approximation that $g_{U}\simeq g_{2}$. To get the correct mass squared, the mass parameter $m_{21}$ plus radiative corrections contribute $\sim$80 GeV to the Higgs Boson mass.
Just as in the Standard Model, it is assumed that radiative corrections of the Higgs Boson are at a scale similar to the electroweak scale, not the unification scale. In the Standard Model, it is an open question as to how cancellation of unification-scale corrections are able to generate electroweak-scale corrections to the Higgs mass. By contrast, supersymmetric theories provide a simple reason for such cancellation: superspace gauge transformations and translations can change a scalar into a fermion, and fermions do not have quadratic divergences. Even if some part of the gauge symmetry is broken at a very large scale in a SUSY theory, as long as the superpotential remains equal to zero, the remaining superspace gauge invariance continues to ensure cancellation of quadratic divergences in the symmetry-broken theory.
Although the theory considered here is not invariant to superspace translations, it is invariant to superspace gauge transformations that change scalars into fermions. Also, the theory’s unification-scale symmetry breaking leaves the superpotential almost equal to zero, with variations from zero only showing up at the electroweak scale. Parallels with supersymmetry then open the door to the possibility that quadratic divergences naturally cancel in this model down to the electroweak scale. Further work would be required to establish this.
@@
On the other hand, interactions of $h_{21}$ with other quarks or leptons are different in this model than in the Standard Model. As described later in this section, leptons and quarks (other than the top) acquire masses from different mechanisms, not from Yukawa interactions with the Higgs Boson. Those particles do have interactions with $h_{21}$ via the nonperturbative terms described in the Appendix. More work would need to be done to determine whether those interactions are consistent with observations.
@@ 3.3 Quark masses and mixing
$\mathit{M}_{U}=\left(\begin{array}[]{ccc}{\hat{g}\bar{\tilde{\phi}}_{21}}&{\tilde{\Gamma}_{211}\bar{\phi}_{21}}&{\tilde{m}_{211}}\\ {\tilde{M}_{G}^{(1)}-\Delta}&{\tilde{m}_{\Phi}}&{\tilde{\Gamma}_{211}\bar{\tilde{\phi}}_{21}}\\ {\tilde{m}_{\lambda}}&{\tilde{M}_{G}^{(1)}+\Delta}&{\hat{g}\bar{\phi}_{21}}\end{array}\right),$
where the gaugino coupling $\hat{g}$ is discussed below in eq (3.14). A tilde is put on tree-level masses and superpotential couplings to show that they include quantum modifications from the Appendix. For all of the fermion mass matrices in this paper (including the above matrix) the rows have an upper undotted spin index (using the convention of [6]) while the columns have a lower undotted spin index (e.g. $\tilde{u}^{\left(1\right)\alpha}u_{W\alpha}^{G}$). Those indices are summed over, and they are suppressed.
@@
The first- and second-generation quark interactions with the $W$ boson, however, differ from those of the Standard Model. For the second-generation up-type quark (mostly charm), both its left- and right-chiral components interact with the $W$ boson. For the first-generation up-type quark (mostly up), only its right-chiral component interacts with the $W$ boson. This chiral flipping of the first generation is not ruled out by experiment, since the spin of the proton (and other light hadrons) comes primarily from gluons and orbital angular momentum [19]. As a result, it is not possible to experimentally disentangle the spins of up and down quarks from other hadronic spins and thereby verify that only left-handed up and down quarks interact with the $W$ boson. Of course, a requirement of this model is that it must be able to reproduce CKM data. It does this even better than the Standard Model, as described below.
@@
Instead, a phenomenological approach is used. Relations for the $W$ boson and top quark masses require
$\hat{g}\simeq\frac{m_{t}}{\sqrt{2}M_{W}}g_{2}\simeq 1.5g_{2}\simeq 0.8g_{3},$ (3.14)
@@
${\rm Rows:}\,\tilde{d}^{(2)},\tilde{d}_{W}^{A},\tilde{d}_{W}^{G},\tilde{d}^{A}\,\,\,{\rm Columns:}\,d_{W}^{G},d^{A},d^{(2)},d_{W}^{A}$ (3.15)
$M_{D}=\left(\begin{array}[]{cccc}{\hat{g}\bar{\tilde{\phi}}_{22}}&{...}&{\tilde{m}_{222}}&{{\tilde{\Gamma}_{222}\bar{\phi}_{22}}}\\ {...}&{\tfrac{1}{\sqrt{2}}\tilde{\Gamma}_{\Phi}c\bar{\varphi}_{2}}&{{\tilde{\Gamma}_{222}\bar{\tilde{\phi}}_{22}}}&{\tilde{m}^{\prime}_{\Phi}}\\ {\tilde{m}^{\prime}_{\lambda}}&{\hat{g}\bar{\varphi}_{2}}&{\hat{g}\bar{\phi}_{22}}&{...}\\ {\hat{g}c\bar{\varphi}_{2}}&{\tilde{m}^{\prime\prime}_{\Phi}}&{...}&{\tfrac{1}{\sqrt{2}}\tilde{\Gamma}_{\Phi}\bar{\varphi}_{2}}\end{array}\right),$
@@
To a first approximation, the masses of the $d$, $s$, and $b$ quarks as well as an additional $f$ quark are given by the diagonal elements of eq (3.15). For example, $m_{b}\simeq\hat{g}\bar{\phi}_{22}$. Comparison with the top quark mass sets the value for $\bar{\phi}_{22}$ at $\bar{\phi}_{22}\simeq\bar{\phi}_{21}m_{b}/m_{t}$. The $f$ quark mass is primarily determined by the adjoint vev $\varphi_{2}$. In this model, none of the down-type quarks get their masses from the Higgs vev $\bar{\phi}_{21}$.
For this model, the numerical values of the quark mass matrices (in GeV) are assumed to be:
${\rm Rows:}\,\bar{u}^{\prime}_{L},\bar{c}^{\prime}_{L},\bar{t}^{\prime}_{L}\,\,\,{\rm Columns:}\,u^{\prime}_{R},c^{\prime}_{R},t^{\prime}_{R}$ (3.16)
$\mathit{M}_{U}=\left(\begin{array}[]{ccc}{-0.0033}&{0.0000}&{-0.0277}\\ {0.1532}&{1.2316}&{0.9688}\\ {7.2510}&{-7.8110}&{171.59}\end{array}\right).$
${\rm Rows:}\,\bar{d}^{\prime}_{L},\bar{s}^{\prime}_{L},\bar{b}^{\prime}_{L},\bar{f}^{\prime}_{L}\,\,\,{\rm Columns:}\,d^{\prime}_{R},s^{\prime}_{R},b^{\prime}_{R},f^{\prime}_{R}$ (3.17)
$\mathit{M}_{D}=\left(\begin{array}[]{cccc}{0.0039}&{-0.0232}&{0.1192}&{0.0503}\\ {-0.0330}&{0.0872}&{0.1953}&{-0.1188}\\ {0.0093}&{-0.0589}&{4.1507}&{0.2934}\\ {0.2497}&{0.0127}&{-0.4219}&{2.8693}\end{array}\right).$
The rows and columns of these matrices are the same gauge eigenstates as in eqs (3.12) and (3.15); they have just been renamed to reflect their dominant mass eigenstate. For example,
$\tilde{d}_{W}^{A}=\bar{s}^{\prime}_{L}=.97\bar{s}_{L}+.22\bar{d}_{L}+.04\bar{f}_{L}-.04\bar{b}_{L}\,,$ (3.18)
where $\bar{s}_{L}$ is in the mass eigenbasis (see below) and $\bar{q}_{L}$ denotes the right-handed field $(q_{L})^{\dagger}$.
The numerical values of these quark mass matrices were chosen by a fit to quark masses and CKM data, while not being in conflict with $Z$, $Z^{\prime}$, and exotic hadron data. In section 5, there is a discussion about how the theory is able to reproduce the many experiments that would seemingly rule out the possibility of an additional light quark. As more data is collected, it may be that rotations of these quark matrices that still reproduce quark masses and CKM data may provide a better global fit. Future work should also verify that there are no inconsistencies between these matrices and the quantum calculations described in the Appendix.
Just as in the Standard Model, each quark mass matrix can be diagonalized via a unitary matrix $V$ on each side:
$V_{L}^{U}\mathit{M}_{U}V_{R}^{U{\dagger}}={\rm diag}(-m_{u},m_{c},m_{t})$ (3.19)
$V_{L}^{D}\mathit{M}_{D}V_{R}^{D{\dagger}}={\rm diag}(-m_{d},m_{s},m_{b},m_{f})\,,$
where the subscript $f$ is used to denote the fourth down-type quark. The unitary matrices that diagonalize the quark mass matrices can be found by first multiplying each mass matrix on the left or on the right by its transpose, then finding the eigenvectors of those product matrices. The allowed flexibility to introduce complex phases was not used for the fits of this paper, so no attempt was made to fit the experimentally measured complex phases of the CKM matrix. The mass eigenvalues of the matrices in eqs (3.16) and (3.17) are:
$m_{u}=0.002,\,\,\,m_{c}=1.28,\,\,\,m_{t}=172$ (3.20)
$m_{d}=0.005,\,\,\,m_{s}=0.095,\,\,\,m_{b}=4.18,\,\,\,m_{f}=2.9\,,$
where all mass values are in GeV.
In the Standard Model, the CKM matrix displays the connections that the $W$ boson makes between up-type and down-type quark mass eigenstates. Since there are 3 quarks of each type in the Standard Model, the CKM matrix is a $3\times 3$ matrix. In this model, there are 3 up-type and 4 down-type quarks, so the equivalent “CKM” matrix is a $3\times 4$ matrix. Also, in the Standard Model only the left-handed quarks (and hence $V_{L}^{U,D}$) matrices have connections to the $W$ boson, but in this model, all four of the matrices have connections to the W.
@@
The placement of the 1’s in the above matrices is based on which quark fields in eqs (3.12) and (3.15) have a $W$ subscript (signifying that they interact with the $W$ boson to change an up-type quark to a down-type quark and vice versa).
In this model, a different CKM matrix should be used depending on whether a vector current or axial vector current process is being considered. Specifically:
@@
Plugging in the unitary matrices $V$ found from eq (3.19) into eq (3.21), the following CKM matrices are obtained:
$\left|V_{\rm{CKM}}^{+}\right|=\left(\begin{array}[]{cccc}{0.9737}&{0.2230}&{0.0039}&{0.0013}\\ {0.2241}&{0.9718}&{0.0394}&{0.9569}\\ {0.0072}&{0.0393}&{0.9934}&{0.0619}\end{array}\right)$ (3.23)
$\left|V_{\rm{CKM}}^{-}\right|=\left(\begin{array}[]{cccc}{0.9732}&{0.2253}&{0.0043}&{0.0012}\\ {0.2183}&{0.9750}&{0.0430}&{1.0414}\\ {0.0824}&{0.0188}&{0.9936}&{0.1451}\end{array}\right).$
@@
In this model, to be consistent with the $W$-boson interactions, the $Z$-boson interactions can be Flavor Changing Neutral Currents (FCNCs). The couplings to the $Z$ boson by quark generation are:
$\begin{array}[]{|c|c|c|}\hline\cr\bar{u}_{i}\gamma^{\mu}u_{i}Z_{\mu}&g_{L}&g_{R}\\ \hline\cr u^{\prime}&-\tfrac{2}{3}x&\tfrac{1}{2}-\tfrac{2}{3}x\\ c^{\prime}&\tfrac{1}{2}-\tfrac{2}{3}x&\tfrac{1}{2}-\tfrac{2}{3}x\\ t^{\prime}&\tfrac{1}{2}-\tfrac{2}{3}x&-\tfrac{2}{3}x\\ \hline\cr{\rm SM}&\tfrac{1}{2}-\tfrac{2}{3}x&-\tfrac{2}{3}x\\ \hline\cr\end{array}$ (3.24)
$\begin{array}[]{|c|c|c|}\hline\cr\bar{d}_{i}\gamma^{\mu}d_{i}Z_{\mu}&g_{L}&g_{R}\\ \hline\cr d^{\prime}&\tfrac{1}{3}x&-\tfrac{1}{2}+\tfrac{1}{3}x\\ s^{\prime}&-\tfrac{1}{2}+\tfrac{1}{3}x&\tfrac{1}{3}x\\ b^{\prime}&-\tfrac{1}{2}+\tfrac{1}{3}x&\tfrac{1}{3}x\\ f^{\prime}&\tfrac{1}{3}x&-\tfrac{1}{2}+\tfrac{1}{3}x\\ \hline\cr{\rm SM}&-\tfrac{1}{2}+\tfrac{1}{3}x&\tfrac{1}{3}x\\ \hline\cr\end{array}\,,$ (3.25)
where $x=\sin^{2}\theta_{W}$ and the “SM” rows show the Standard Model couplings for up-type and down-type quarks. The primed quark variables are gauge eigenstates, just as in eqs (3.16) and (3.17).
Since the couplings of this model differ by generation, $Z$ boson connections mix mass eigenstates. For example, the Z mixing matrix for left-chiral down-type quarks is:
$V_{ZD}^{L}$ $=V_{L}^{D}\left(\begin{array}[]{cccc}{\tfrac{1}{3}x}&{0}&{0}&{0}\\ {0}&{-\tfrac{1}{2}+\tfrac{1}{3}x}&{0}&{0}\\ {0}&{0}&{-\tfrac{1}{2}+\tfrac{1}{3}x}&{0}\\ {0}&{0}&{0}&{\tfrac{1}{3}x}\end{array}\right)V_{L}^{D{\dagger}}$ (3.26)
$=\left(\begin{array}[]{cccc}{0.0520}&{-0.1082}&{0.0141}&{0.0066}\\ {-0.1082}&{-0.3970}&{-0.0056}&{0.0221}\\ {0.0141}&{-0.0056}&{-0.4172}&{-0.0505}\\ {0.0066}&{0.0221}&{-0.0505}&{0.0709}\end{array}\right),$
where the diagonalizing matrix $V_{L}^{D}$ is the same as was used for CKM matrices, and $x=\sin^{2}\theta_{W}=0.2315$ was used. There are also $Z$ mixing matrices for right-chiral down-type quarks as well as for right- and left-chiral up-type quarks.
Most of the off-diagonal FCNC elements of these matrices are small, but a few are more significant. The only Z mixing matrix elements with magnitudes greater than 0.006 are $V_{Zs\bar{d}}^{A},V_{Zf\bar{b}}^{L},V_{Zf\bar{s}}^{L},V_{Zf\bar{d}}^{L},V_{Zb\bar{d}}^{L}$, $V_{Zt\bar{c}}^{R}$, and $V_{Zt\bar{u}}^{R}$. More detailed work would be needed to verify that the $Z$ boson FCNC mixing of this model is fully consistent with experimental data.
3.4 Leptons
In this model, charged leptons get their masses not from the vev of the Higgs boson but from vevs of the $\varphi_{2}$ adjoint scalars (see eq (3.4)). Through the superpotential couplings of eq (1.18), $m_{1FF^{\prime}}$ and $\Gamma_{1FF^{\prime}}$ generate charged lepton masses. Combining those with quantum-generated masses from (A.12) leads to the following lepton mass terms: